OpenStudy (anonymous):

Match each expression with its solution. Match Term Definition three hundred divided by ten times three minus the quantity one plus two raised to the third power, end quantity A) ninety-seven three hundred divided by the quantity ten times three, end quantity, minus one plus two raised to the third power B) sixty-eight three hundred divided by ten times the quantity three minus one, end quantity, plus two raised to the third power C) eighty-one three hundred divided by ten times three minus one plus two raised to the third power D) seventeen

OpenStudy (anonymous):

Just read the sentence carefully and try to make expressions out of it.. :)

OpenStudy (anonymous):

Wait, what is quantity here??

OpenStudy (anonymous):

I got it..

OpenStudy (anonymous):

Look for the second statement here.. the quantity ten times three will mean : \((10 \times 3)\) and only ten times three means : \(10 \times 3\) (I mean no brackets)..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, let us try to decode Second Statement: three hundred means 300. Divided by means : / the quantity ten times 3 means : \((10 \times 3)\) minus 1 means : -1 Plus means : + 2 raised to power third means : \((2)^3\)

OpenStudy (anonymous):

300 / 30 -1* 8 right ???

OpenStudy (anonymous):

So, here use those statements to make your expression as : 1. 300 2. 300/ 3. \(300 \div(10 \times 3)\) 4. \(300 \div (10 \times 3) - 1\) 5. \(300 \div (10 \times 3) -1 + (2)^3\)

OpenStudy (anonymous):

So, finally for second statement : \[\frac{300}{10 \times 3} - 1 + (2)^3 \implies \frac{300}{30} - 1 + 8 \implies \color{green}{17}\]

OpenStudy (anonymous):

Are you getting it??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Note that, the whole quantity ie (10*3) is getting divided by 300..

OpenStudy (anonymous):

Third is also same, so we will first try that..

OpenStudy (anonymous):

It is simple for you to understand that: the quantity means : Brackets open there... end quantity means : Brackets close there. Like in third statement, the quantity three minus one, end quantity means : (3 - 1)

OpenStudy (anonymous):

Are you getting this "the quantity" and "end quantity" concept??

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

So, if you are getting, then look at first statement and try to use the same quantity concept there and try to tell me what you will get there inside the brackets??

OpenStudy (anonymous):

Third Statement will decode to: \[\frac{300}{10} \times (3-1) + (2)^2 \implies 30 \times 2 + 8 \implies \color{green}{68}\]

OpenStudy (anonymous):

300/30 - (9)= 1 right ????

OpenStudy (anonymous):

For which statement you have written it??

OpenStudy (anonymous):

First statement